# Joint-Embedding Masked Autoencoder for Self-supervised Learning of Dynamic Functional Connectivity from the Human Brain

### v1

@@ IV-B Block Masking Process for Masked Autoencoding
As presented in Assran et al. [2], when training a model with a JEPA-based masked autoencoding objective, a sufficiently large size of the target masks, with sequential indices, is crucial in the image dataset. This strategy ensures that the masked targets encompass meaningful semantic information, allowing models to grasp the semantic structure by reconstructing the target data from the remaining context. Although this can be easily done by masking the patch for the image data, applying similar masking techniques to nodes and edges in graph data is challenging due to the absence of clear relationships between sequential indices for nodes. However, in the context of dynamic FC of fMRI data, the graph construction method outlined in Kim et al. [17] defines a well-aligned construction between nodes with adjacent indices across time. Consequently, masking nodes with sequential indices is a natural and effective approach for the dynamic FC of fMRI data. This method allows for the masking of nodes without sacrificing meaningful semantic information.
As mentioned earlier, generating adequately sized masks for both nodes and edges at each time step $t$ is essential to capture meaningful semantic information within the masked blocks. To achieve this for node features $\boldsymbol{X}(t)$, we generate $K$ binary mask matrices $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)$, where each $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)$ randomly and independently masks a $\lfloor\alpha_{\texttt{node}}^{(k)}(t)\cdot N\rfloor$ number of nodes with sequential indices. Here, mask ratio $\alpha_{\texttt{node}}^{(k)}(t)$ uniformly sampled from $(\alpha_{\text{min}},\alpha_{\text{max}})$ where $10<\alpha_{\text{min}}<\alpha_{\text{max}}<$ 1 are predefined hyperparameters. This process ensures the creation of appropriately sized blocks. Refer to Appendix Section E-B for the experimental results regarding the ablation study on block mask ratios.
@@
$\boldsymbol{X}_{\texttt{cxt}}(t)=\boldsymbol{X}(t)\odot\bigg{(}\boldsymbol{M}^{(1)}_{\boldsymbol{X}}(t)\odot\ldots\odot\boldsymbol{M}^{(K)}_{\boldsymbol{X}}(t)\bigg{)},$ (12)
where $\odot$ denotes elementwise product.
@@
$\boldsymbol{A}_{\texttt{cxt}}(t)=\boldsymbol{A}(t)\odot\bigg{(}\boldsymbol{M}^{(1)}_{\boldsymbol{A}}(t)\odot\ldots\odot\boldsymbol{M}^{(K)}_{\boldsymbol{A}}(t)\bigg{)}.$ (13)
@@
IV-C Spatial Reconstruction Process
Based on the masks generated from Section IV-B, we introduce an objective function inspired by JEPA, referred to as ${\mathcal{L}}_{\texttt{spat}}$, aiming to facilitate the learning of spatial patterns within the dynamic FC of fMRI data using GNN representations. To create ${\mathcal{L}}_{\texttt{spat}}$, we need to reconstruct the target node representations and the target adjacency matrix based on the context node representations and the context adjacency matrix. To achieve this reconstruction, we initially build the common context node representation for each time $t$ as follows:
@@ IV-C Spatial Reconstruction Process
$\begin{gathered}\tilde{\boldsymbol{Z}}(t)=\boldsymbol{W}_{\texttt{S}}^{{\mathcal{V}}}\boldsymbol{Z}(t),\\ \tilde{\boldsymbol{Z}}_{\texttt{cxt}}(t)=\tilde{\boldsymbol{Z}}(t)\odot\Big{(}\boldsymbol{M}_{\boldsymbol{X}}^{(1)}(t)\odot\cdots\odot\boldsymbol{M}_{\boldsymbol{X}}^{(K)}(t)\Big{)},\end{gathered}$ (17)
where $\boldsymbol{W}_{\texttt{S}}^{{\mathcal{V}}}\in\mathbb{R}^{d_{\texttt{dec}}\times d_{\texttt{enc}}}$ denotes a learnable projection matrix that is used to adjust from the encoder dimension $d_{\texttt{enc}}$ to the decoder dimension $d_{\texttt{dec}}$. Using the masked context node representations $\tilde{\boldsymbol{Z}}(t)$ and any node mask $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)$, we derive the target-specific context node representations $\tilde{\boldsymbol{Z}}_{\texttt{cxt}}^{(k)}(t)$ as follows:
@@
where ${\mathcal{L}}_{\texttt{MSE}}$ denotes the Mean Squared Error (MSE) loss. In contrast to prior graph-based MAE methods [28, 14], where the entire node representations are recovered by the decoder module, our approach only focuses on recovering the target node representations. This strategy guides the model to concentrate on learning spatial patterns by specifically emphasizing the retrieval of semantic information embedded within the masked target nodes.
@@
IV-D Temporal Reconstruction Process
To enhance the understanding of temporal dynamics within the dynamic FC of fMRI data, we also construct a JEPA-based ${\mathcal{L}}_{\texttt{temp}}$. This temporal reconstruction strategy is designed to predict node representations and adjacency matrices at a specific time $t$ by utilizing node representation from two random points in time, $t_{a}$ and $t_{b}$, that can encompass $t$ within their interval. For the selection of $t_{a}$ and $t_{b}$, we adopt a uniform sampling strategy following Choi et al. [7], a straightforward method for choosing two temporal points. An important consideration when sampling $t_{a}$ and $t_{b}$ is the use of a temporal windowing process with a length of $\Gamma$ and a stride of $S$ when constructing a series of FC matrices from the BOLD signal, as discussed in Section III-C. If $t_{a}$ and $t_{b}$ are sampled too closely, there may be an overlap between the windows of $t_{a}$ and $t_{b}$, making the reconstruction task trivial. To ensure the challenge and effectiveness of the temporal reconstruction task, we take care to ensure that the sampled time points $t_{a}$ and $t_{b}$ do not originate from overlapping windows as follows:
@@ IV-D Temporal Reconstruction Process
Employing $\boldsymbol{Z}_{a,b}(t)$ to populate the target node representations encourages our model to better grasp the temporal dynamics within the data, enhancing its understanding of temporal variations. Then our final recovered target node representations $\bar{\boldsymbol{Z}}_{\texttt{tar}}^{(k)}(t)$ is computed as below:
@@
Here, we used the same decoder $g_{\texttt{node}}$ within the spatial reconstruction. Our temporal node reconstruction loss ${\mathcal{L}}_{\texttt{node-temp}}^{(t)}$ at time $t$ is computed as:
@@
IV-E Summary of the Overall ST-JEMA Pipeline
@@
where the hyperparameter $\gamma\in[0,1]$ is the ratio of ${\mathcal{L}}_{\texttt{spat}}$. The overview and complete training pipeline of the Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA) are described in Fig. 1 and Appendix Appendix A.
V Experimental Setup
We utilize the UKB [27] as the upstream dataset, a large-scale fMRI data comprised of 40,913 subjects, for pre-training GNN encoder with SSL. We evaluate the pre-trained GNN encoder across eight downstream datasets both non-clinical and clinical rs-fMRI datasets. For non-clinical datasets, we evaluate gender classification and age regression tasks using ABCD  [5], and three types of HCP datasets: HCP-YA [33], HCP-A [3], and HCP-D [26]. For clinical datasets, we extended psychiatric diagnosis classification tasks to distinguish between normal control groups and different psychiatric disorder groups using HCP-EP [20], ABIDE [9], ADHD200 [4], and COBRE [23].
@@ V Experimental Setup
The training process was divided into two phases: 1) pre-training, where a GNN encoder $f(;{\boldsymbol{\theta}})$ and decoder $g(;{\boldsymbol{\phi}})$ were learned using SSL on the UKB dataset without phenotype labels, employing a 4-layer GIN [35] for encoder module and a one-layer MLP-mixer for the decoder in the ST-JEMA framework, and 2) fine-tuning, where the pre-trained encoder and a task-specific head $h(;{\boldsymbol{\omega}})$, incorporating the SERO readout module[17], were further trained on downstream tasks with supervised learning. Pre-training employed a cosine decay learning schedule over 10K steps with a batch size of 16. Fine-tuning optimized the encoder with the task-specific head, using orthogonal regularization and a one-cycle scheduler for 30 epochs with a batch size of 32. Dynamic graphs integrated a mean pooling module for final representation processing. Classification tasks used cross-entropy loss, while age regression utilized MSE loss with normalized age data. The training utilized AMD EPYC processors, with NVIDIA RTX A6000 for pre-training and GeForce RTX 3090 for fine-tuning and ablation studies.
For further details on the experimental setup, please refer to Appendix Appendix B, Appendix C, and Appendix D.
VI Results
TABLE I: Gender classification performance across various downstream fMRI datasets. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
@@ VI Results
TABLE II: Age regression performance across various downstream fMRI datasets. Performance is evaluated using the MAE score, with lower scores indicating better performance.
@@
TABLE III: Psychiatric diagnosis classification performance on clinical fMRI datasets. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
@@
TABLE IV: Linear probing performance of psychiatric diagnosis classification. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
Pre-training AUROC ($\uparrow$) HCP-EP ABIDE ADHD200 COBRE ST-DGI 70.66 53.87 53.32 63.73 ST-MAE 70.78 55.09 52.66 63.97 ST-JEMA $\mathbf{72.97}$ $\mathbf{58.30}$ $\mathbf{54.44}$ $\mathbf{66.39}$
TABLE V: Multi-task learning performance on ABIDE and ADHD200 datasets. Here, (-) represents a degree of performance drop, and (+) indicates the degree of performance improvement compared to the performance in each single-task learning scenario.
Pre-training ABIDE Gender Age Diagnosis AUROC ($\uparrow$) MAE ($\downarrow$) AUROC ($\uparrow$) ST-DGI 65.00 (-9.87) 4.24 (-0.08) 59.50 (-9.36) ST-MAE 62.86 (-11.88) 4.11 (-0.01) 58.25 (-11.48) ST-JEMA $\mathbf{65.60}$ (-9.78) $\mathbf{4.06}$ (+0.02) $\mathbf{66.10}$ (-5.39) Pre-training ADHD200 Gender Age Diagnosis AUROC ($\uparrow$) MAE ($\downarrow$) AUROC ($\uparrow$) ST-DGI 60.06 (-7.98) 1.99 (-0.08) 55.61 (-0.85) ST-MAE 60.98 (-10.03) $\mathbf{1.89}$ (-0.03) 50.45 (-6.32) ST-JEMA $\mathbf{68.29}$ (-3.65) $1.91$ (-0.05) $\mathbf{59.84}$ (+0.95)
VI-A Downstream-task Performance
For gender and psychiatric disorder classification, we measured the Area Under the Receiver Operating Characteristic (AUROC) curve, while for the age regression task, we utilized the MAE as a measure of prediction accuracy. In all tables, the best results are in bold, second-best results are $\underline{\text{underlined}}$, and ‘Rank’ shows average rankings across datasets/tasks.
VI-A1 Gender Classification
@@
In gender classification tasks using the gender class phenotype from benchmark datasets, our proposed method, ST-JEMA, demonstrates superior AUROC performance across all benchmarks as shown in Table I. Notably, it proved especially effective in classifying gender from clinical datasets, which typically comprise fewer samples than non-clinical datasets. We noted that employing dynamic rather than static FC resulted in performance improvements when comparing the overall performance of static and dynamic methods. This outcome proves the significance of considering temporal variations in gender classification tasks. Additionally, ST-JEMA outperformed previous dynamic graph self-supervised learning methods. This enhancement in performance shows that ST-JEMA has effectively learned representations that capture the temporal dynamics in the upstream rs-fMRI data, facilitated by our proposed loss function ${\mathcal{L}}_{\texttt{ST-JEMA}}$. This improvement enhances its applicability for phenotype classification.
VI-A2 Age Regression
@@
In the age regression task, ST-JEMA consistently surpassed other methods in prediction accuracy across the most of datasets. The evaluation metric scales vary due to different age ranges in each dataset, refer to Table II. Interestingly, the performance of the GNN encoder trained via ST-DGI was similar to, or in some instances worse than, the Dynamic Random-Init method, especially in clinical datasets that have fewer samples. This indicates that the representations learned through ST-DGI, a contrastive-based SSL method, may not consistently contribute positively to learning information in downstream tasks. On the other hand, ST-JEMA showed improved performance even with clinical data, suggesting that the representations learned through our methodology were more effective at distinguishing phenotypes.
VI-A3 Psychiatric Diagnosis Classification
@@
The motivation behind employing SSL with a large amount of unlabeled fMRI data is to improve the performance of the downstream diagnosis classification task, typically constrained by a limited number of labeled training datasets. We aim for the GNN encoder to first understand the structure of fMRI data from upstream learning, and then, with minimal clinical data, focus on distinguishing psychiatric differences for improved classification. As shown in Table III, in most cases, SSL-trained model showed better performance than those initialized randomly, proving that learning representations from unlabeled data is beneficial for accurately detecting signal variations in BOLD signals for psychiatric diagnosis classification. ST-JEMA notably surpassed other methods in psychiatric diagnosis classification, especially demonstrating significantly higher AUROC in the COBRE dataset, even when tasked with distinguishing between three distinct psychiatric conditions and control groups.
VI-B Linear Probing Performance
According to the results in Table IV, ST-JEMA exhibited higher AUROC performance than ST-MAE, suggesting that ST-JEMA captures more semantically rich information beneficial for psychiatric diagnosis classification. On the contrary, ST-DGI, a SSL method based on contrastive learning, exhibited lower performance than generative approaches. This suggests that the representations learned by ST-DGI might not be as beneficial for psychiatric diagnosis classification tasks compared to generative-based SSL methods.
VI-C Multi-task Learning Performance
Table V clearly shows that ST-JEMA achieved the best performance across all tasks compared to other baselines except for the age regression task in the ADHD200 dataset. Notably, ST-JEMA exhibits the smallest overall performance decline for each task when compared to the performance in the single-task learning scenario, outperforming other baseline models. In particular, regarding the age regression task in the ABIDE dataset and diagnosis classification task in ADHD200, ST-JEMA demonstrates an improvement in performance in the multi-task learning scenario compared to its performance in the single-task learning scenario. These findings highlight that the pre-trained encoder in ST-JEMA has effectively captured mutual information beneficial for addressing diverse downstream tasks.
VI-D Ablation Study

### v2

@@ 4.2 Block Masking Process for Masked Autoencoding
As presented in Assran et al. [14], when training a model with a JEPA-based masked autoencoding objective, a sufficiently large size of the target masks, with sequential indices, is crucial in the image dataset. This strategy ensures that the masked targets encompass meaningful semantic information, allowing models to grasp the semantic structure by reconstructing the target data from the remaining context. While patch-based masking is straightforward for images, applying similar masking techniques to nodes and edges in graph data is challenging due to the absence of clear relationships between sequential indices for nodes. However, in the context of dynamic FC of fMRI data, the graph construction method outlined in Kim et al. [5] defines a well-aligned construction between nodes with adjacent indices across time. Consequently, masking nodes with sequential indices is a natural and effective approach for the dynamic FC of fMRI data. This method allows for the masking of nodes without sacrificing meaningful semantic information.
As mentioned earlier, generating adequately sized masks for both nodes and edges at each time step $t$ is essential to capture meaningful semantic information within the masked blocks. To achieve this for node features $\boldsymbol{X}(t)$, we generate $K$ binary mask matrices $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)$, where each $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)$ randomly and independently masks a $\lfloor\alpha_{\texttt{node}}^{(k)}(t)\cdot N\rfloor$ number of nodes with sequential indices. Here, mask ratio $\alpha_{\texttt{node}}^{(k)}(t)$ uniformly sampled from $(\alpha_{\text{min}},\alpha_{\text{max}})$ where $10<\alpha_{\text{min}}<\alpha_{\text{max}}<$ 1 are predefined hyperparameters. This process ensures the creation of appropriately sized blocks. Refer to Section F.5 for the experimental results regarding the ablation study on block mask ratios.
@@
$\boldsymbol{X}_{\texttt{cxt}}(t)=\boldsymbol{X}(t)\odot\bigcap_{k=1}^{K}\boldsymbol{M}^{(k)}_{\boldsymbol{X}}(t),$ (7)
where $\odot$ denotes element-wise (Hadamard) product and $\bigcap$ denotes element-wise logical AND over binary matrices.
@@
$\boldsymbol{A}_{\texttt{cxt}}(t)=\boldsymbol{A}(t)\odot\bigcap_{k=1}^{K}\boldsymbol{M}^{(k)}_{\boldsymbol{A}}(t),$ (8)
@@
4.3 Spatial Reconstruction Process
Based on the masks generated from Section 4.2, we introduce an objective function inspired by JEPA, referred to as ${\mathcal{L}}_{\texttt{spat}}$, aiming to facilitate the learning of spatial patterns within the dynamic FC of fMRI data using GNN representations. To create ${\mathcal{L}}_{\texttt{spat}}$, we need to reconstruct the target node representations and the target adjacency matrix based on the context node representations and the context adjacency matrix. To achieve this reconstruction, we initially build the common context node representation for each time $t$ as follows:
@@ 4.3 Spatial Reconstruction Process
$\begin{gathered}\tilde{\boldsymbol{Z}}(t)=\boldsymbol{W}_{\texttt{S}}^{{\mathcal{V}}}\boldsymbol{Z}(t),\\ \tilde{\boldsymbol{Z}}_{\texttt{cxt}}(t)=\tilde{\boldsymbol{Z}}(t)\odot\bigcap_{k=1}^{K}\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t),\end{gathered}$ (12)
where $\boldsymbol{W}_{\texttt{S}}^{{\mathcal{V}}}\in\mathbb{R}^{d_{\texttt{dec}}\times d_{\texttt{enc}}}$ projects encoder outputs to decoder space. Using the masked context node representations $\tilde{\boldsymbol{Z}}(t)$ and any node mask $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)$, we derive the target-specific context node representations $\tilde{\boldsymbol{Z}}_{\texttt{cxt}}^{(k)}(t)$ as follows:
@@
where ${\mathcal{L}}_{\texttt{MSE}}$ denotes the Mean Squared Error (MSE) loss. In contrast to prior graph-based MAE methods [27, 8], where the entire node representations are recovered by the decoder module, our approach only focuses on recovering the target node representations. This encourages the model to focus on semantic information embedded within the masked target nodes.
@@
4.4 Temporal Reconstruction Process
To enhance the understanding of temporal dynamics within the dynamic FC of fMRI data, we also construct a JEPA-based ${\mathcal{L}}_{\texttt{temp}}$. This temporal reconstruction strategy is designed to predict node representations and adjacency matrices at a specific time $t$ by utilizing node representation from two random points in time, $t_{a}$ and $t_{b}$, that can encompass $t$ within their interval. For the selection of $t_{a}$ and $t_{b}$, we adopt a uniform sampling strategy following Choi et al. [13], a straightforward method for choosing two temporal points. An important consideration when sampling $t_{a}$ and $t_{b}$ is the use of a temporal windowing process with a length of $\Gamma$ and a stride of $S$ when constructing a series of FC matrices from the BOLD signal, as discussed in Appendix B. If $t_{a}$ and $t_{b}$ are sampled too closely, there may be an overlap between the windows of $t_{a}$ and $t_{b}$, making the reconstruction task trivial. To ensure the challenge and effectiveness of the temporal reconstruction task, we take care to ensure that the sampled time points $t_{a}$ and $t_{b}$ do not originate from overlapping windows as follows:
@@ 4.4 Temporal Reconstruction Process
Employing $\boldsymbol{Z}_{a,b}(t)$ to populate the target node representations helps the model capture temporal dynamics more effectively, enhancing its understanding of temporal variations. The recovered target node representations are obtained using the same decoder $g_{\texttt{node}}$ as in the spatial task:
@@
Our temporal node reconstruction loss ${\mathcal{L}}_{\texttt{node-temp}}^{(t)}$ at time $t$ is computed as:
@@
4.5 Summary of the Overall ST-JEMA Pipeline
@@
where the hyperparameter $\gamma\in[0,1]$ is the ratio of ${\mathcal{L}}_{\texttt{spat}}$. The overview and complete training pipeline of the ST-JEMA are described in Fig. 1 and Appendix A.
5 Experimental Setup
We utilize the UK Biobank [28] as the upstream dataset, a large-scale fMRI data comprised of 40 913 subjects, for pre-training GNN encoder with SSL. We evaluate the pre-trained GNN encoder across eight downstream datasets both non-clinical and clinical rs-fMRI datasets. For non-clinical datasets, we evaluate gender classification and age regression tasks using ABCD  [29], and three types of HCP datasets: HCP-YA [30], HCP-A [31], and HCP-D [32]. For clinical datasets, we extended psychiatric diagnosis classification tasks to distinguish between normal control groups and different psychiatric disorder groups using HCP-EP [33], ABIDE [34], ADHD200 [35], and COBRE [36].
@@ 5 Experimental Setup
The training process was divided into two phases: 1) pre-training, where a GNN encoder $f(;{\boldsymbol{\theta}})$ and decoder $g(;{\boldsymbol{\phi}})$ were learned using SSL on the UKB dataset without phenotype labels, employing a 4-layer GIN [38] for encoder module and a one-layer MLP-mixer for the decoder in the ST-JEMA framework, and 2) fine-tuning, where the pre-trained encoder and a task-specific head $h(;{\boldsymbol{\omega}})$, incorporating the SERO readout module[5], were further trained on downstream tasks with supervised learning.
Further details on the experimental setup are provided in the dataset specifications (Appendix C), baseline descriptions (Appendix D), and training configurations (Appendix E).
6 Results
Table 1: Gender classification performance across various downstream fMRI datasets. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
@@ 6 Results
Table 2: Age regression performance across various downstream fMRI datasets. Performance is evaluated using the MAE score, with lower scores indicating better performance.
@@
Table 3: Psychiatric diagnosis classification performance on clinical fMRI datasets. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
@@
6.1 Downstream-task Performance
6.1.1 Gender Classification
@@
In gender classification tasks using the gender class phenotype from benchmark datasets, our proposed method, ST-JEMA, demonstrates superior AUROC performance across all benchmarks as shown in Table 1. Notably, it proved especially effective in classifying gender from clinical datasets, which typically comprise fewer samples than non-clinical datasets. We noted that employing dynamic rather than static FC resulted in performance improvements when comparing the overall performance of static and dynamic methods. This outcome proves the significance of considering temporal variations in gender classification tasks. Additionally, ST-JEMA outperformed previous dynamic graph self-supervised learning methods. This enhancement in performance shows that ST-JEMA has effectively learned representations that capture the temporal dynamics in the upstream rs-fMRI data.
6.1.2 Age Regression
@@
In the age regression task, ST-JEMA consistently surpassed other methods in prediction accuracy across the most of datasets. The evaluation metric scales vary due to different age ranges in each dataset, refer to Table 2. Interestingly, the performance of the GNN encoder trained via ST-DGI was similar to, or in some instances worse than, the Dynamic Random-Init method, especially in clinical datasets that have fewer samples. Contrastive methods like ST-DGI showed unstable results, while ST-JEMA showed improved performance even in clinical data, suggesting that the representations learned through our methodology were more effective in distinguishing phenotypes.
6.1.3 Psychiatric Diagnosis Classification
@@
The motivation behind employing SSL with a large amount of unlabeled fMRI data is to improve the performance of the downstream diagnosis classification task, typically constrained by a limited number of labeled training datasets. The GNN encoder is expected to learn the structure of fMRI data during upstream training, enabling better psychiatric classification with minimal clinical data. As shown in Table 3, in most cases, SSL-trained model showed better performance than those initialized randomly, proving that learning representations from unlabeled data is beneficial for accurately detecting signal variations in BOLD signals for psychiatric diagnosis classification. ST-JEMA notably surpassed other methods in psychiatric diagnosis classification, especially demonstrating significantly higher AUROC in the COBRE dataset, even when tasked with distinguishing between three distinct psychiatric conditions and control groups.
We further evaluated the versatility of representations learned by ST-JEMA through linear probing, multi-task learning, and temporal missing data scenarios. These results support the robustness of our model across diverse settings. See Appendix F for details.
6.2 Ablation Study